OpenStudy (anonymous):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 9 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section?

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

Hey

OpenStudy (qwertty123):

P(red) = 11/20 P(blue) = 6/20 P(green) = x

OpenStudy (qwertty123):

P(red) + P(blue) + P(green) = 1 11/20 + 6/20 + x = 1 solve for x

OpenStudy (qwertty123):

brb Hope this makes sense! I need to help someone else.

OpenStudy (qwertty123):

what is 11/20 + 6/20 equal to

OpenStudy (qwertty123):

@loveyouthanks

OpenStudy (qwertty123):

17/20

OpenStudy (qwertty123):

so we have 17/20 + x = 1

OpenStudy (qwertty123):

subtract 17/20 from both sides

OpenStudy (qwertty123):

x = 1 - (17/20) x = 20/20 - 17/20 ... think of 1 as 20/20 x = ???